OpenStudy (rockindud):

The mass of P4O10 that will be obtained from the reaction of 1.33 gram of P4 and 5.07 of oxygen is

OpenStudy (sohailiftikhar):

ok here first see that for one molecule of p4O10 how much mass of p4 and O10 is required

OpenStudy (sohailiftikhar):

p=31g O=16g P4=3(31) O=10(16) P4O10=93+160 =253g

OpenStudy (sohailiftikhar):

this is for one molecule

OpenStudy (sohailiftikhar):

now it looks like that we are given the mass of P4 9 times smaller than the orignal =93/9=1.33g

OpenStudy (sohailiftikhar):

so their should be 9 times smaller mass of O10 is required to form a product ..so in this case product will not form

OpenStudy (sohailiftikhar):

what are the options you have ?

OpenStudy (sohailiftikhar):

????

OpenStudy (rockindud):

[A]. 2.05 gram [B]. 3.05 gram @ [C]. 4.05 gram [D]. 5.05 gram